OpenStudy (anonymous):

A car traveled 960 mi at a certain speed. If the speed had been 12 miles faster, the trip could have been made in 4 hr less time. Find the speed.

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

not at all...

OpenStudy (anonymous):

wow!! if "not at all", then "how to start"?? hihihi

OpenStudy (anonymous):

I don't know that's why im asking for help:(

OpenStudy (anonymous):

Let get help @mosaic

OpenStudy (anonymous):

can't think of a snappy way, but i bet we can do it

OpenStudy (anonymous):

Yayy.. Thank you!!

OpenStudy (anonymous):

gimme a second to see if i can come up with an easy way elsewise we just grind it til we find it

OpenStudy (anonymous):

ok

OpenStudy (mosaic):

Let x be the speed. 960 / x - 960 / (x+12) = 4

OpenStudy (anonymous):

distance is rate times time, so one equation is \[960=RT\] and the other is \[960=(R+12)(T-4)\]

OpenStudy (anonymous):

@mosaic might be easier i was going to suggest \[T=\frac{960}{R}\] and solve \[(R+12)(\frac{960}{R}-4)=960\]

OpenStudy (anonymous):

probably the same either way

OpenStudy (anonymous):

@KaylaAnn!79 you got a preference?

OpenStudy (anonymous):

Not really but the second one looks promising

OpenStudy (anonymous):

k then we have to multiply all this mess out

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

wanna cheat?

OpenStudy (anonymous):

looks like the answer is \(48\) but if you want we can do all the stupid algebra to get it